### v2

# Long-Range Angular Correlations of Particle Displacements at a Plastic-to-Elastic Transition in Jammed Amorphous Solids

Abstract
Understanding how a fluid turns into an amorphous solid is a fundamental challenge in statistical physics, during which no apparent structural ordering appears. In the athermal limit, the two states are connected by a well-defined jamming transition, near which the solid is marginally stable. A recent mechanical response screening theory proposes an additional transition above jamming, called a plastic-to-elastic transition here, separating anomalous and quasielastic mechanical behavior. Through numerical inflation simulations in two dimensions, we show that the onsets of long-range radial and angular correlations of particle displacements decouple, occurring, respectively, at the jamming and plastic-to-elastic transitions. The latter is characterized by a power-law diverging correlation angle and a power-law spectrum of the displacements along a circle. This work establishes two-step transitions on the mechanical properties during “decompression melting” of an athermal overjammed amorphous solid, reminiscent of the two-step structural melting of a crystal in two dimensions. In contradistinction with the latter, the plastic-to-elastic transition exists also in three dimensions.
preprint: APS/123-QED
Introduction. When strained by an increasing deformation $\gamma$, a crystal displays a crossover from elastic to plastic responses. If the crystal is compressed, it responds elastically to a small $\gamma$ at any pressure $p$. The picture is dramatically changed in athermal amorphous solids near the jamming transition ($p\approx 0$). Such a solid is known to be marginally stable [1, 2, 3], with elasticity breaking down in the thermodynamic limit, even to infinitesimal mechanical perturbations [4, 5]. Indeed, the minimum strain $\gamma_{\rm min}$ required to trigger a plastic event vanishes in large systems [6, 7]. On the other hand, one expects elasticity restored in overjammed solids well above the jamming transition ($p\gg 0$), when the interparticle overlapping $\delta$ exceeds expected $\gamma_{\rm min}$, such that under small deformations the system can be effectively considered as an elastic medium. Thus there should be a plastic-to-elastic (PE) transition (or crossover) moving away from jamming.
The expected PE transition has been recently explored by two approaches, both of which originated from theories. The first approach builds on the mean-field replica theory (MFRT) [3, 8], which predicts a Gardner transition [3, 9, 10, 11, 12] in athermal soft spheres at $\varphi_{\rm G}^{>}$ above the jamming density $\varphi_{\rm J}$ [5]. The Gardner transition separates the elastic phase and the plastic phase where the shear modulus is protocol dependent due to marginal stability [13, 14, 15, 16].
Figure 1: Radial displacement fields. Each row corresponds to typical response behavior at different pressures (from top to bottom): $p=0,4.4\times 10^{-2},5.2\times 10^{2},8.6\times 10^{2}$, for $N=67352$. First column: heat maps of $\hat{d}_{r}(x,y)$ normalized by the maximum $\hat{d}_{r}^{\rm max}$. Second column: $\hat{d}_{r}(r)$, where the lines in (h,k) represent Eq. (1) with $d_{0}^{\rm eff}=0.054$ and 0.048, respectively, and the line in (e) represents Eq. (2) with $d_{0}^{\rm eff}=0.072$ and $\kappa_{\rm fit}=0.026$. Third column: $\hat{d}_{r}(\theta)$ at $\hat{r}\equiv r/r_{\rm out}=0.15,0.5$, and $0.85$ (purple, green, and red, respectively).
The second approach is motivated by a recent mechanical response screening theory (MRST) [17, 18, 19, 20, 21, 22]. The theory provides a possible PE transition in athermal overjammed solids, between a quasielastic phase with quadrupole screening of the elastic field by plastic events, where plasticity simply renormalizes the elastic moduli, and an anomalous phase with dipole screening. This transition is observed in recent simulations, by showing the change of the screening parameter $\kappa$ from zero to a finite value at a nonzero pressure $p_{\rm c}$ [23]. However, several crucial questions remain to be answered regarding the nature of the transition: in particular, whether a sharp transition or a crossover occurs at $p_{\rm c}$ in the thermodynamic limit, and in the former case, whether it is associated with diverging long-range correlations.
Following the second approach, here we show that the PE transition can be unambiguously differentiated from the jamming transition: the jamming transition is associated with a diverging length due to long-range radial correlations in displacements, and, in contrast, the PE transition corresponds to a diverging angle due to long-range transverse correlations. It should be noted that the PE transition captures the change of the macroscopic mechanical response behavior, which is beyond the first contact change events at the microscopic level [24, 25].
Three phases: quasielastic, plastic and unjammed. We simulate a bidisperse granular model of Hertzian (unless otherwise specified) particles in two dimensions, confined between a fixed outer boundary of radius $r_{\rm out}$ and an inner boundary of radius $r_{\rm in}$. A small particle, with a unit diameter ($R_{1}=0.5$) and a fixed position at the center ($r=0$), is inflated as $r_{\rm in}\to r_{\rm in}+d_{0}$, where $r_{\rm in}=R_{1}=0.5$ and $d_{0}=0.1$ (unless otherwise specified). The energy of the system is minimized after this instantaneous inflation, and the particle responses are characterized by the displacement field $\bm{\mathrm{d}}(r,\theta)$. In this study, we focus on the radial component of the displacement $d_{r}(r,\theta)$, where $d_{r}=\bm{\mathrm{d}}\cdot\hat{\bm{\mathrm{n}}}$ and $\hat{\bm{\mathrm{n}}}$ is the unit radial vector. The transverse component $d_{\theta}(r,\theta)$ is investigated in a separate study [26], revealing an interesting odd dipole screening effect. Details of the model and simulation methods are described in Appendixes A and B.
The mechanical responses to the above-described inflation can be categorized into three phases:
(i) A quasielastic solid phase at $p>p_{\rm c}$ and $\varphi>\varphi_{\rm c}$, where the mechanical response in the radial direction is elasticlike. The displacement field is dominated by dipolar and quadrupolar patterns [see Figs. 1(g)-1(l)]. Although the central inflation is isotropic, the actual displacement and force fields are generally anisotropic. Hexapoles and higher-order poles exist but are originated from $1/f^{1.5}$ noise (see Fig. 3 and related discussions).
The displacement field $d_{r}(r,\theta)$ in this phase is consistent with the Michell solution that describes the mechanical response of a standard elastic medium [27]. Using the rescaled radial displacements $\hat{d}_{r}(r,\theta)=rd_{r}(r,\theta)$ to remove the trivial $1/r$ decay of $d_{r}(r,\theta)$ at large $r$ and averaging out the angular dependence, the radial dependence of $\hat{d}_{r}(r)$ is expected to follow the solution of the normal elasticity theory,
$\hat{d}_{r}(r)=d_{0}\frac{r_{\rm in}(r^{2}-r_{\rm out}^{2})}{(r_{\rm in}^{2}-r_{\rm out}^{2})}.$ (1)
Eq. (1) is consistent with simulation results, with the inflation variable treated as a fitting parameter $d_{0}\to d_{0}^{\rm eff}$ [see Figs. 1(h) and 1(k)].
Eq. (1) suggests that the response is elasticlike. The application of inflation may result in plastic events, which are typically quadrupolar in nature, aka Eshelby inclusions. At a large $p$, the plastic events are sparse, and the gradient of the quadrupole field $\bm{\mathrm{Q}}(r,\theta)$ is negligible. In this case, the effect of plasticity is to renormalize the elastic moduli, without modifying the overall behavior of the displacement field Eq. (1). For this reason, the high-pressure phase is termed a quasielastic phase. However, the angular dependence of $\hat{d}_{r}$ is not isotropiclike in a true elastic medium (see Sec. S1 in Supplemental Material (SM) [28]). The oscillative behavior of $\hat{d}_{r}(\theta)$ (at a given $r$) suggests long-range correlations in the angular direction [Figs. 1(i) and 1(l)).
(ii) A plastic solid phase at $0<p<p_{\rm c}$ and $\varphi_{\rm J}<\varphi<\varphi_{\rm c}$, where the system is overall jammed but the mechanical response anomalously disobeys elasticity. The displacement field $\hat{d}_{r}(r,\theta)$ can not be described by any regular patterns [see Figs. 1(d)-(f)]. Under the isotropic assumption, the screening theory provides a solution of $\hat{d}_{r}(r)$ [17],
$\hat{d}_{r}(r)=d_{0}r\frac{Y_{1}(r\kappa)J_{1}(r_{\rm out}\kappa)-J_{1}(r\kappa)Y_{1}(r_{\rm out}\kappa)}{Y_{1}(r_{\rm in}\kappa)J_{1}(r_{\rm out}\kappa)-J_{1}(r_{\rm in}\kappa)Y_{1}(r_{\rm out}\kappa)},$ (2)
where $J_{1}$ and $Y_{1}$ are circular Bessel functions of the first and second kind, respectively, and $\kappa$ is a screening parameter. Simulation data are fitted to Eq. (2) with two fitting parameters, $d_{0}^{\rm eff}$ and $\kappa_{\rm fit}$ [see Fig. 1(e)]. Note that Eq. (1) is equivalent to setting $\kappa=0$ in Eq. (2). See Sec. S2 in SM [28, 29, 30] for ensemble-averaged data.
The screening parameter $\kappa\sim 1/\ell_{\rm s}$ is the inverse of the screening length $\ell_{\rm s}$. The screening effect originates from an effective dipole field $\bm{\mathrm{P}}(r,\theta)$, which is the gradient of $\bm{\mathrm{Q}}(r,\theta)$ when the quadrupolar events are dense and non-uniformly distributed in the space [19]. While Eq. (2) gives rise to long-range correlations in the radial direction, $d_{r}(\theta)$ is noiselike, implying short-range correlations in the angular direction [Fig. 1(f)].
(iii) An unjammed phase at $p=0$ and $\varphi<\varphi_{\rm J}$, where the system is in a fluid state. The influence of the central inflation propagates only up to a finite distance $\xi$, and $d_{r}(\theta)$ is noiselike [see Figs. 1(a)-1(c)].
Figure 2: Plastic-to-elastic transition. (a) The screening parameter $\kappa_{\rm fit}$ from fitting, and (b) the scaled pressure difference $\Delta p/p^{1/3}$ as functions of $p$, for $N=67352$. The blue data points (averaged over samples with $\kappa_{\rm fit}<0.01$ at the given $p$; error bars are invisible) have near-zero values, which can be unambiguously distinguished from others (red points; each point represents one sample).
Let us highlight the essential differences between the three phases by summarizing the radial and angular dependence of $\hat{d}_{r}$ (see Fig. 1): (i) in the quasielastic solid phase, $\hat{d}_{r}(r)$ follows the elasticity theory Eq. (1) and $\hat{d}_{r}(\theta)$ is oscillative (both long ranged); (ii) in the plastic solid phase, $\hat{d}_{r}(r)$ follows the MRST Eq. (2) and $\hat{d}_{r}(\theta)$ is noiselike (long ranged in $r$ and short ranged in $\theta$); (iii) in the unjammed phase, $\hat{d}_{r}(r)$ decays to zero at a finite $\xi$ and $\hat{d}_{r}(\theta)$ is noiselike (both short ranged).
Two transitions: jamming and plastic-to-elastic. The jamming transition between unjammed and plastic phases has been extensively investigated in previous studies [31, 32, 33, 34, 35, 36, 1, 37]. In Ref. [34], the radial correlations between particles’ nonaffine displacements are measured in athermal quasistatic shear simulations with Lees-Edwards boundary conditions, giving a diverging correlation length $\xi$ when the jamming transition is approached from below,
$\xi\sim(\varphi_{\rm J}-\varphi)^{-\nu},$ (3)
with $\nu\sim 0.8-1.0$. In this study, we estimate $\xi$ by the length scale at which the radial displacement $d_{r}(r)$ decays to a small threshold value, $d_{r}(r=\xi)=10^{-3}$ (see Sec. S3 in SM [28]). The power-law divergence of $\xi$ is confirmed with $\varphi_{\rm J}=0.842$ [38] and $\nu\approx 0.8$ [34] consistent with previous studies.
The PE transition between plastic and quasielastic phases was only noticed very recently [23]. To provide a rough estimation of the PE transition pressure $p_{\rm c}$, we follow the strategy in  [23]: the data of $\hat{d}_{r}(r)$ at different $p$ are fitted to Eq. (2), and the fitting parameter $\kappa_{\rm fit}$ is plotted as a function of $p$ in Fig. 2(a). At high $p$, $\kappa_{\rm fit}\approx 0$ (blue points), consistent with Eq. (1). The nonzero $\kappa_{\rm fit}$ appears only below $p_{\rm c}\approx 20$ for the system of $N=67352$ particles (red points). In Ref. [23], a selection principle of $\kappa$ is proposed: the $\hat{d}_{r}(r)$ curves measured in simulations preferably select a series of $\kappa_{n}\approx n\pi/r_{\rm out}$ with a decreasing probability when $n$ is increased, where $n=1,2,\ldots$ At $\kappa=\kappa_{n}$ the screening effects are significantly strong, which can be seen from the behavior of Eq. (2), since the amplitude $\hat{d}_{\rm amp}(\kappa=\kappa_{n})$ diverges. Here $\hat{d}_{\rm amp}(\kappa)$ is the extreme value of the $\hat{d}_{r}(r)$ curve for the given $\kappa$. In Fig. 2(a), the first two $\kappa_{n}$, i.e., $\kappa_{1}\approx 0.022$ and $\kappa_{2}\approx 0.041$, are indicated by dashed lines. The data points of nonzero $\kappa_{\rm fit}$ all fall in the range between $\kappa_{1}$ and $\kappa_{2}$. The clear gap between $\kappa=0$ and $\kappa_{1}$ supports the selection principle. Meanwhile, we observe a wide fluctuation of $\kappa_{\rm fit}$ at $p<p_{\rm c}$, possibly due to strong sample-to-sample fluctuations in finite-sized systems and the uncertainty in the fitting.
Figure 3: Angular correlation functions and power spectrum. (a) Angular correlation function $G(\theta)$ and (b) power spectrum $S(f)$ at different $\hat{p}=pN^{3/2}$, for $\hat{r}=r/r_{\rm out}=0.5$ and $N=67352$. The solid lines in (b) represent fitting to Eq. (5).
To examine the reliability of $p_{\rm c}$ estimated above, in this study we propose two additional methods to determine $p_{\rm c}$. In Fig. 2(b), we plot the pressure difference $\Delta p\equiv p_{\rm final}-p$ between the pressures after and before the central inflation (both pressures are measured with the energy minimized). To properly normalize the pressure, note that $p\sim F\sim\delta^{3/2}$, where the second relation between the interparticle force $F$ and the interparticle linear overlap $\delta$ holds for the Hertzian potential. From this, one obtains $d\delta\sim dp/p^{1/3}$. The data (red points) in Fig. 2(b) shows that $\Delta p/p^{1/3}$ fluctuates around $-0.05$ below $p_{\rm c}\approx 20$. The fluctuation is large, but nevertheless the negative values confirm the anomaly, compared to the small positive values expected from the elastic behavior (blue points). Note that this second approach can be straightforwardly accessed in granular experiments [18]. In the next section, we will discuss a third way to determine $p_{\rm c}$, which gives the most accurate estimate.
Figure 4: Order parameters and correlation angles. $G^{*}$ and $\theta^{*}$ as functions of (a,b) $p$ and (c,d) $\hat{p}$, for $\hat{r}=0.5$ and different $N$. Open symbols in (d) are $\theta^{\dagger}$ data. (inset) $\theta^{\dagger}$ as a function of $(\hat{p}_{\rm c}-\hat{p})/\hat{p}_{\rm c}$, with the power-law fitting to Eq. (6).
Emergence of long-range angular correlations at the plastic-to-elastic transition. As discussed above, the onset of long-range displacement correlations in the radial direction occurs at the jamming transition $p=p_{\rm J}=0$ and $\varphi=\varphi_{\rm J}$. Next we show that the onset of long-range displacement correlations in the angular direction appears at the PE transition $p=p_{\rm c}$. We define an angular correlation function of the radial displacements,
$G(\theta)=\overline{c\langle\delta{d}_{r}(\theta_{0})\,\delta{d}_{r}(\theta_{0}+\theta)\rangle},$ (4)
[…]
Figure  3(a) reports $G(\theta)$ for $\hat{r}=0.5$ at different pressures. In the large pressure limit, the displacement field is dominated by dipolar patterns as shown in Fig. 1(j), which corresponds to $G(\theta)=\cos(\theta)$. With decreasing $p$, quadrupolar and other patterns [see Fig. 1(g)] appear, and $G(\theta)$ deviates from the cosine function. Below $p_{\rm c}$, $G(\theta)$ rapidly decays to zero, indicating short-range correlations. To quantify the behavior change of $G(\theta)$, we define an order parameter $G^{*}\equiv-G(\theta=\pi)$, and a correlation angle $G(\theta=\theta^{*})=G_{\rm th}$ with a threshold $G_{\rm th}=0.1$ (see Sec. S4 in SM [28]). Figures  4(a) and 4(b) show that both $G^{*}$ and $\theta^{*}$ grow with $p$.
The data of $G^{*}$ and $\theta^{*}$ suffer strong finite-size effects and thus it is necessary to consider a proper finite-size scaling. We find that they can be collapsed when plotted as functions of a rescaled pressure, $\hat{p}=pN^{3/2}$ [Figs. 4(c) and  4(d)]. To understand this finite-size scaling, recall that near the jamming transition, the isostatic length $\ell_{\rm iso}\sim(Z-Z_{\rm iso})^{-1}\sim p^{-1/3}$ sets a natural length scale, where $Z$ is the average coordination number and $Z_{\rm iso}=2d=4$ is the isostatic coordination number in $d=2$ dimensions [39]. According to the argument in Ref. [23], the PE transition occurs when $\ell_{\rm iso}$ matches the largest screening length $\ell_{\rm s}\sim 1/\kappa_{1}\approx r_{\rm out}/\pi\sim N^{1/2}$, which gives $p_{\rm c}\sim N^{-3/2}$, consistent with our data (see SM  [28]). For a more general interaction potential $U(\delta)\sim\delta^{\alpha}$, $p_{\rm c}(N,d_{0})$ of different $N$ and inflation $d_{0}$ obeys a scaling function $N^{\beta}p_{\rm c}(N,d_{0})=\mathcal{G}(N^{2}d_{0})$, where $\beta=2(\alpha-1)$ and $\mathcal{G}(x)\sim x^{\beta/4}$ (see Appendix D and Sec. S5 in SM [28]).
Next we demonstrate that the system has long-range angular correlations at $\hat{p}_{\rm c}$, even though $\theta^{*}$ does not appear to diverge. In our setup, the long-range correlation corresponds to the convergence, $\theta^{*}\to\theta^{*}_{\rm c}\approx 0.3\pi$ and $G^{*}\to G^{*}_{\rm c}\approx 0.3$, when $\hat{p}\to\hat{p}_{\rm c}$ from below. To show that, we perform the Fourier transform of $d_{r}(\theta)$ and plot the (normalized) power spectrum $S(f)$ in Fig. 3(b) at different $\hat{p}$, for $N=67352$. Interestingly, $S(f)$ displays a power-law with an exponent $-1.5$ (which we term $1/f^{1.5}$ noise) at $\hat{p}_{\rm c}\approx 10^{8}$, indicating angular scale-free behavior. The determined $\hat{p}_{\rm c}\approx 10^{8}$ is consistent with $\hat{p}_{\rm c}\approx 3\times 10^{8}$ estimated in Fig. 2 (considering strong data fluctuations there) and $\hat{p}_{\rm c}\approx 8\times 10^{7}$ reported in Ref. [23] (converted from $p_{\rm c}\approx 3.5$ for $N=80000$). In the analyses below we take $\hat{p}_{\rm c}\approx 10^{8}$ as the value of the rescaled transition pressure. In Sec. S6 of SM [28], we show that the $1/f^{1.5}$ behavior at $\hat{p}_{\rm c}$ is robust against varying $\hat{r}$, and is universal in Hertzian and Hookean systems.
With the power-law behavior at $\hat{p}_{\rm c}$ established, it becomes possible to characterize the growth of angular correlations from the spectrum data. We fit the $S(f)$ data at $\hat{p}<\hat{p}_{\rm c}$ to a generalized Lorentzian form,
$S_{\rm GL}(f)=\frac{A}{1+(f/f^{\dagger})^{1.5}},$ (5)
where $f^{\dagger}$ is the only fitting parameter, and $A=1+(1/f^{\dagger})^{1.5}$ is fixed by $f^{\dagger}$ due to the normalization $S(f=1)=1$. Here $f^{\dagger}$ separates the white noise at $f<f^{\dagger}$ with $S(f)\sim 1$ from the $1/f^{1.5}$ noise at $f>f^{\dagger}$. Since $f^{\dagger}$ is a characteristic wave number, one can define a corresponding angle, $\theta^{\dagger}\equiv k/f^{\dagger}$, with $k=0.45$ estimated in Sec. S7 of SM [28, 40]. The physical meaning of $\theta^{\dagger}$ can be interpreted as follows: the angular correlation between displacements exists up to $\theta^{\dagger}$, beyond which the displacements are uncorrelated like white noise. Assuming uncorrelated displacements in the fluid state, one expects that $\theta^{\dagger}\to 0$ approaching the jamming transition. On the other hand, the scale-free spectrum $S(f)$ at $\hat{p}_{\rm c}$ suggests that $\theta^{\dagger}\to\infty$ (i.e., $f^{\dagger}\to 0$) approaching the PE transition. Moreover, in the plastic phase, $S(f)$ deviates from the scale-free scaling due to a finite $\theta^{\dagger}$. These expectations are consistent with the data in Fig. 3(b). In particular, near $\hat{p}_{\rm c}$, $\theta^{\dagger}$ diverges such that
$\theta^{\dagger}\sim(\hat{p}_{\rm c}-\hat{p})^{-\mu},$ (6)
where $\mu\approx 0.9$ from fitting [see Fig. 4(d) (inset)]. Eq. (6) is the angular counterpart of Eq. (3). Note that $\theta^{\dagger}$ diverges at $p_{\rm c}$ (or $\varphi_{\rm c}$) and $\xi$ diverges at $p_{\rm J}$ (or $\varphi_{\rm J}$).
The two correlation angles, $\theta^{*}$ and $\theta^{\dagger}$, coincide at small $\hat{p}$ [Fig. 4(d)]. The angle $\theta^{*}$ does not diverge at $p_{\rm c}$, because it is defined in a finite domain $\theta\in[0,\pi]$. On the other hand, $\theta^{\dagger}$ can diverge due to the analytic extension of the frequency, $f^{\dagger}=k/\theta^{\dagger}\to 0$. In linear inflation geometry, we expect that $S(f)\sim f^{-1.5}$ would still hold at $p_{\rm c}$ and the associated length scale perpendicular to the inflation direction would diverge without the need of analytic extension.
Discussion. The angular behavior of the particle displacements $d_{r}(\theta)$ in response to the central inflation can be summarized as follows. (i) In the unjammed phase ($p=0$), the displacements are uncorrelated and white-noiselike, with a zero correlation angle. (ii) In the plastic phase ($0<p<p_{\rm c}$), the correlation angle is finite, and the displacements are correlated only within the correlation angle. (iii) In the quasielastic phase ($p>p_{\rm c}$), the correlation angle is finite because the displacement field is dominated by dipolar and quadrupolar patterns. The long-range angular correlation emerges at the PE transition ($p_{\rm c}$), captured by a diverging correlation angle $\theta^{\dagger}$ and a power-law spectrum $S(f)\sim f^{-1.5}$.
This study opens several questions for future studies. In particular, the origin of the $f^{-1.5}$ spectrum at $p_{\rm c}$ needs to be recovered. It will be useful to generalize the current setup of circular inflation to other types of deformations [41, 42, 43]. Despite the similarity between the current problem and two-dimensional melting, unlike the latter, we expect that the two-step transitions discussed here would also appear in higher dimensions [21]. Finally, it would be very interesting to reconcile the MFRT and MRST approaches.
Acknowledgments. We warmly thank Matteo Baggioli, Bulbul Chakraborty, Wouter Ellenbroek, Wenyue Fan, Michael Moshe, Jin Shang, and Jie Zhang for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 12047503 and 12404290), from Chinese Academy of Sciences (Grant ZDBS-LY-7017), and from Wenzhou Institute (Grant WIUCASQD2023009). IP acknowledges support from the ISF under grant #3492/21 (collaboration with China) and the Minerva Center for “Aging, from physical materials to human tissues” at the Weizmann Institute. In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
Supplementary Information
S1 Dipolar and quadrupolar patterns of the displacement fields at $p>p_{\rm c}$ under square periodic boundary conditions
In the quasi-elastic regime ($p>p_{\rm c}$), dipolar and quadrupolar patterns of the displacement fields are commonly observed (see Fig. 1(g,j)). To examine if such anisotropic patterns are caused by the fixed circular boundary conditions, we perform additional inflation simulations in a square simulation box with periodic boundary conditions (the center of mass of the entire system is fixed). As shown in Fig. S1, dipolar and quadrupolar patterns are repeatedly observed in different samples, confirming that the anisotropic patterns at $p>p_{\rm c}$ do not significantly depend on the boundary conditions.
Figure S1: Radial displacement fields under periodic boundary conditions. Presented in (a-j) are the heap maps of $\hat{d}_{r}(x,y)$ obtained for ten independent samples of $N=2774$ disks at $p=3.5\times 10^{3}$.
S2 Ensemble-averaged radial displacements
The data of rescaled radial displacements $\hat{d}_{r}(r)$ presented in Fig. 1(b,e,h,k) are obtained for single configurations. The single-sample curves are suitable to fit the solutions of the mechanical response screening theory (MRST). Previous inflation simulations have also considered ensemble-averaged quantities, which are averaged over many independent configuration samples [29, 30]. The ensemble-averaged $\overline{\hat{d}_{r}(r)}$ are presented in Fig. S2, for a few different pressures $p$ and inflation parameters $d_{0}$. For the largest $p$ and smallest $d_{0}$, the behavior of $\overline{\hat{d}_{r}(r)}$ is close to the elastic solution Eq. (1). With decreasing $p$ or increasing $d_{0}$, deviations from Eq. (1) are observed. When $p$ is near or below $p_{\rm c}$, or when $d_{0}$ is near or larger than $d_{\rm c}$, the ensemble-averaged $\overline{\hat{d}_{r}(r)}$ mix samples with elastic and anomalous responses, and thus they can be fitted by neither Eq. (1) nor Eq. (3). In order to compare with theoretical solutions (Eq. (1) or (3)), one should use single-sample data $\hat{d}_{r}(r)$ instead of the ensemble-averaged data $\overline{\hat{d}_{r}(r)}$. Fig. S2 also shows that the coarse-grained elasticity  [29, 30] works in the quasi-elastic phase ($p>p_{\rm c}$ and $d_{0}<d_{\rm c}$), but fails in the plastic phase ($p<p_{\rm c}$ and $d_{0}>d_{\rm c}$) where nonlinear behavior is important.
Figure S2: Ensemble-averaged $\overline{\hat{d}_{r}(r)}$ for $N=67352$. (a) A few different pressures $p$, for a fixed $d_{0}=0.1$. (b) A few different $d_{0}$, for a fixed $p=9\times 10^{-1}$.
S3 Jamming transition
The unjammed configurations with $p=0$ are studied and it is found that the displacement response of the central inflation propagates only up to a finite distance. Without loss of generality, we define the length scale $\xi$ by $d_{r}(r=\xi)=d_{\rm th}=10^{-3}$. The power-law divergence Eq. (3) of $\xi$ with $\nu\approx 0.8$ is confirmed in Fig. S3 when approaching $\varphi_{\rm J}=0.842$, consistent with previous studies  [38, 34]. The exponent $\nu$ is unchanged when the threshold $d_{\rm th}$ in the definition of $\xi$ is decreased from $10^{-3}$ to $10^{-6}$ (see Fig. S4).
Figure S3: Jamming transition. The characteristic length $\xi$ for the jamming transition as a function of (a) the packing density $\varphi$, and (b) $\varphi_{\rm J}-\varphi$, for a few different $N$.
Figure S4: Correlation length $\xi$ obtained using different thresholds. The threshold $d_{\rm th}$ in the definition $d_{r}(r=\xi)=d_{\rm th}$ is respectively (a) $10^{-4}$, (b) $10^{-5}$ and (c) $10^{-6}$. The exponent $\nu\approx 0.8$ is unchanged.
S4 Dependence of $\theta^{*}$ on the threshold $G_{\rm th}$
Figure S5 shows that the finite-size scaling $p_{\rm c}\sim N^{-3/2}$ in Hertzian systems is nearly independent of the threshold value $G_{\rm th}$ used in the definition of $\theta^{\ast}$, $G(\theta=\theta^{\ast})=G_{\rm th}$.
Figure S5: Correlation angle $\theta^{*}$ as a function of $pN^{3/2}$, using different threshold values $G_{\rm th}$. In (a-c), $G_{\rm th}=0.02,0.05,0.2$ respectively.
S5 Additional data on the scaling analysis in Hookean systems
In Hertzian systems, we have shown the $N$-dependence of $p_{\rm c}$, $p_{\rm c}\sim N^{-3/2}$, in the main text. For Hookean disks, the numerical results suggest a finite-size scaling $p\sim N^{-1}$, instead of $p\sim N^{-3/2}$, for a fixed $d_{0}$ (see Fig. S6). In addition, the data of $\theta^{*}$ and $G^{*}$ can be collapsed as functions of $pN/d_{0}^{1/2}$ (see Fig. S7).
Figure S6: Finite-size scaling in Hookean disks. The simulation data of $G^{*}$ and $\theta^{*}$ can be collapsed as functions of (a,b) $pN$, but not (c,d) $pN^{3/2}$. The inflation $d_{0}=0.1$ is fixed.
Figure S7: Dependence of the PE transition on the inflation $d_{0}$ and the system size $N$, for Hookean systems. (a) $G^{*}$ and (b) $\theta^{*}$ as functions of $pN/d_{0}^{1/2}$, for three different $N$ and four different $d_{0}$, where the dotted line marks $p_{\rm c}N/d_{0}^{1/2}$. (c) The $N$-rescaled $d_{0}-p$ phase diagram, where the dotted line is $(N^{2}d_{0})\sim(N^{2}p_{\rm c})^{2}$.
S6 Dependence of the power spectrum on the radius
Fig. S8 shows that the power-law behavior of the power spectrum $S(f)\sim f^{-1.5}$ at $\hat{p}_{\rm c}$ is robust against varying $\hat{r}$, although the boundary effects inevitably kink in when $r$ is close to $r_{\rm out}$. Note that the data collapse for $\hat{r}<0.6$. The exponent $1.5$ is observed in both Hertzian and Hookean potentials.
Figure S8: Power spectrum $S(f)$ at the PE transition $\hat{p}_{\rm c}$ for (a) Hertzian and (b) Hookean potentials, with $N=67352$. The dotted line represent a power-law decay $f^{-1.5}$.
S7 Inverse Fourier transform of the generalized Lorentzian spectrum
According to the Weiner-Khintchine theorem [40], the power spectrum $S(f)$ of $d_{r}(\theta)$ and its auto-correlation $G(\theta)$ form a Fourier transform pair. The inverse Fourier transform of the generalized Lorentzian spectrum (Eq. (5) in the main text) gives,
$G_{\rm GL}(\theta)=\sum_{f=1}^{\infty}\frac{B\cos(f\theta)}{1+(f/f^{\dagger})^{1.5}},$ (S1)
where $B=\sum_{f}\frac{1}{1+(f/f^{\dagger})^{1.5}}$ is the normalizing factor, and $f$ is an integer because $d_{r}(\theta)$ is periodic in $\theta\in[0,2\pi]$. Following the main text, we define $\theta^{*}_{\rm GL}$ by $G_{\rm GL}(\theta=\theta^{*}_{\rm GL})=0.1$. As shown in Fig. S9, the curve $\theta^{*}_{\rm GL}(f^{\dagger})$ is close to the simulation data $\theta^{*}$ (for $\hat{p}<\hat{p}_{\rm c}$). In the large $f^{\dagger}$ limit (i.e., $\hat{p}\to 0$), $\theta^{*}_{\rm GL}(f^{\dagger}\to\infty)=k/f^{\dagger}$, with $k\approx 0.45$. Based on this, we define $\theta^{\dagger}=k/f^{\dagger}$ in the main text. In the small $f^{\dagger}$ limit (i.e., $\hat{p}\to\hat{p}_{\rm c}$), $\theta^{*}_{\rm GL}(f^{\dagger}\to 0)=\theta_{\rm GL,c}^{*}\approx 0.3\pi$. In addition, $G^{*}_{\rm GL,c}\equiv-G_{\rm GL}(\theta=\pi)\approx 0.3$.
According to the above analyses, we conclude that $\theta^{*}$ and $G^{*}$ converge to finite values at $\hat{p}_{\rm c}$, $\theta^{*}_{\rm c}\approx 0.3\pi$ and $G^{*}_{\rm c}\approx 0.3$. In other words, even though $S(f)\sim f^{-1.5}$ is clearly scale-free at $\hat{p}_{\rm c}$, its inverse Fourier transform $G(\theta)$ does not show a diverging correlation angle because it has to satisfy the periodic boundary conditions at $\theta=0$ and $\theta=2\pi$, such that $f$ needs to be an integer with minimum $f=1$ in Eq. (S1). This constraint comes from the current setup of circular inflation. If one instead performs line inflation, then the real variable ($f=n/L$) analytic extension is possible for Fourier modes $\cos\left(\frac{2n\pi x}{L}\right)$, when the linear size of the system $L\to\infty$.
Figure S9: Correlation angle from the inverse Fourier transform of the generalized Lorentzian spectrum. $\theta_{\rm GL}^{*}$ as a function of $f^{\dagger}$ (red line). Data points are $\theta^{*}$ obtained from simulations (we only show $\theta^{*}$ with $\hat{p}<\hat{p}_{\rm c}$).
S8 Isostatic length
The isostatic length $\ell_{\rm iso}$ has been determined explicitly in previous central inflation simulations, using the fluctuation data of the change in contact force, $\Delta F$ [29, 30]. For a pair of contacting particles $i$ and $j$, by definition the force change $\Delta F_{ij}$ is proportional to $d_{\parallel,ij}$, i.e., $\Delta F_{ij}\equiv-k_{n}d_{\parallel,ij}$, where $d_{\parallel,ij}$ is the parallel component of the relative displacement $\bm{\mathrm{d}}_{ij}=\bm{\mathrm{d}}_{j}-\bm{\mathrm{d}}_{i}$ along the contact bond, and $k_{n}$ is the spring coefficient. Inspired by this method, here we use the fluctuations of $d_{\parallel}$ to extract $\ell_{\rm iso}$. Note the difference between $d_{r}$ and $d_{\parallel}$: $d_{r}$ is the projection of the displacement $\bm{\mathrm{d}}$ onto the radial direction with respect to the inflation center of the simulation system; $d_{\parallel}$ is the projection of the relative displacement $\bm{\mathrm{d}}_{ij}$ onto the bond direction, which represents the change in the bond length. Comparing the visualized fields of $d_{\parallel}(r,\theta)$ in Fig. S10 with the corresponding $d_{r}(r,\theta)$ in Fig. 1(d, g, j), one sees that $d_{\parallel}(r,\theta)$ captures the physics of isostatic stability, rather than that of the plastic screening effects. The $d_{\parallel}(r,\theta)$ field looks disordered within a length scale $\ell_{\rm iso}$ from the inflation center. At larger length scales beyond $\ell_{\rm iso}$, the role of disorder is smeared out and the $d_{\parallel}(r,\theta)$ field looks more similar to what is expected for a continuous elastic medium. In other words, $d_{r}(r,\theta)$ and $d_{\parallel}(r,\theta)$ encode respectively the information of the screening length $\ell_{\rm s}$ and the isostatic length $\ell_{\rm iso}$.
In order to evaluate $\ell_{\rm iso}$ quantitatively, we follow Refs. [29, 30] and compute the root mean square fluctuations $h_{\parallel}(r)=\sqrt{\left\langle\left(d_{\parallel}(r)-\langle d_{\parallel}(r)\rangle\right)^{2}\right\rangle}$, where the average $\langle x\rangle$ is taken over along the circle of radius $r$. As shown in Fig. S11, the data of $h_{\parallel}(r)$ at different $p$ all collapse when plotted as functions of $r\Delta Z$ with $\Delta Z=Z-Z_{\rm iso}$, suggesting the characteristic length scale $\ell_{\rm iso}\sim 1/\Delta Z$. The tail of $h_{\parallel}(r)$ follows a power-law $h_{\parallel}(r)\sim r^{-1}$, corresponding to the regime where the effects of disorder disappear. The threshold to reach this asymptotic power-law defines a proper pre-factor for the scaling of $\ell_{\rm iso}$ (see Fig. S11b),
$\ell_{\rm iso}=\frac{C}{\Delta Z},$ (S2)
with $C\approx 6$, consistent with Refs. [29, 30]. We have checked that the fluctuations of $d_{r}(r)$ do not collapse as functions of $r\Delta Z$ (see Fig. S11(c,d)).
Figure S10: Visualization of the $d_{\parallel}$ field at (a) $p=4.4\times 10^{-2}$, (b) $p=5.2\times 10^{2}$ and (c) $p=8.6\times 10^{2}$, for $N=67452$ and $d_{0}=0.1$. The red (blue) lines indicate compressed (stretched) bonds with $d_{\parallel}<0$ ($d_{\parallel}>0$). The black circles indicate $\ell_{\rm iso}$.
Figure S11: Fluctuations of $d_{\parallel}$. (a) The root mean square fluctuations $h_{\parallel}(r)$ as functions of $r\Delta Z$ for different $p$. The deviations from the master curve at large $r$ are due to boundary effects. (b) $h_{\parallel}(r)\times(r\Delta Z)$ as functions of $r\Delta Z$, showing a plateau at large distances. We choose a threshold $h_{\parallel}(r)\times(r\Delta Z)=1.3$ (dashed line) that gives the crossover value $\ell_{\rm iso}\Delta Z=C\approx 6$. For comparison, the fluctuations of $d_{r}(r)$, defined by $h_{r}(r)=\sqrt{\left\langle\left(d_{r}(r)-\langle d_{r}(r)\rangle\right)^{2}\right\rangle}$ are plotted in (c,d), which do not collapse as functions of $r\Delta Z$.
S9 Contact changes
Figure S12 shows that the inflation of the central particle causes multiple contact changes in our simulations. Thus the plastic-to-elastic (PE) transition does not correspond to the first contact change event analyzed in previous studies [24, 25].
Figure S12: Contact changes due to (a-c) inflation and (d-f) deflation of the central disk. The pressures are (a,d) $p=4.4\times 10^{-2}$, (b,e) $p=5.2\times 10^{2}$, (c,f) $p=8.6\times 10^{2}$ (corresponding to the second to fourth rows in Fig. 1). The inflation-deflation simulations are performed following the procedure $d_{0}=0\to 0.1\to 0$, and the configurations are equilibrated after the inflation and deflation. The contact networks before and after the inflation/deflation are compared. The broken and created contacts are marked by blue and red points respectively. Note that there is no contact change for unjammed systems since $Z=0$.